OpenStudy (anonymous):

Find the equation of the circle with center at (3, 2) and through the point (5, 4).

OpenStudy (anonymous):

I dont need the answer just a little help?

OpenStudy (shaik0124):

With a center at point (-1, -2), the formula will be (x+1)^2 + (y+2)^2 = r^2. As the point mentioned is on the circle, the distance from the canter point to this point must be the radius r. With the point (15,-2) on the circle, the distance from this point to the center is 16. The radius is therefore 16. The formula is therefore: (x+1)^2 + (y+2)^2 = 16^2 or (x+1)^2 + (y+2)^2 = 256

OpenStudy (anonymous):

Thank you!! I understand it better now.

OpenStudy (shaik0124):

ok :)